OpenStudy (anonymous):

H11p1 formulae anyone?

OpenStudy (anonymous):

H11P1: a) 1/(RC*C) + RL/L b) 56.5 c) d) 14.6 e) H11P2 a) Vm*0.1 = 0.05 b) 1.44 c) 1.44 d) VM(t)*0.1 = 0.8 H11P3 a) C b) R c) L d) 4 e) L f) 1000 g) 1 h) 16

OpenStudy (anonymous):

Q3 - range of ans is within 4 and 10 Q5 - range of ans is within 15 and 20 also check for more nearer values if not getting correct or else gimme the values of F in both d question :)

OpenStudy (anonymous):

hello :-)

OpenStudy (anonymous):

H11P1 1.ω=2πf Capacitance C=1/((ω^2)L) 2.Bandwidth Δω = RL/L + 1/(Rc *C) 3.Capacitance C=1/((ω^2)L) 4.Bandwidth Δω = RL/L + 1/(Rc *C)